OpenStudy (anonymous):

how to find the number of x-intercepts in the graph of the equation y=3x2 – 3x + 4

OpenStudy (rajee_sam):

x intercepts are the points where the graph cuts the x axis. at those points values of y=0. So it is just another way of asking you to find the zeros aka solve for x

Parth (parthkohli):

Yes, so you have to solve the equation \(3x^2 - 3x + 4 = 0\)

OpenStudy (anonymous):

3(x-4)(x-1)=0 ; divide by 3? ; x=5 and x=-1 ???

OpenStudy (anonymous):

3(x-4)(x+1)=0 ; divide by 3? ; x=4 and x=-1 ??? typo